OpenStudy (anonymous):

Find the inverse of the quadratic function. y=x^2 -1

OpenStudy (bibby):

Switch x and y and solve for y(well y'). x=\(y^2-1\)

OpenStudy (morganchavez):

f^-1(x)=\[\sqrt{x+1},-\sqrt{x+1}\]

OpenStudy (morganchavez):

that is the inverse of the function

OpenStudy (anonymous):

oh dang you guys beat me XD

OpenStudy (morganchavez):

lol

OpenStudy (anonymous):

So +-\[\sqrt{x+1}\]?

OpenStudy (bibby):

\(y'=\pm \sqrt{x+1}\)

OpenStudy (anonymous):

ok ty